# Reducing quantum error correction overhead using soft information

### v2

@@ Abstract
Imperfect measurements are a prevalent source of error across quantum computing platforms, significantly degrading the logical error rates achievable on current hardware. To mitigate this issue, rich measurement data referred to as soft information has been proposed to efficiently identify and correct measurement errors as they occur. In this work, we model soft information decoding across a variety of physical qubit platforms and decoders and showcase how soft information can make error correction viable at lower code distances and higher physical error rates than is otherwise possible. We simulate the effects of soft information decoding on quantum memories for surface codes and bivariate bicycle codes, and evaluate the error suppression performance of soft decoders against traditional decoders. For near-term devices with noise regimes close to the surface code threshold, our simulations show that soft information decoding can provide up to $10\%$ higher error suppression on superconducting qubits and up to $20\%$ stronger error suppression on neutral atom qubits.
@@ I Introduction
Improvements in classical decoders can offer a shortcut to stronger error suppression, but the accuracy of the decoder is often limited by how much information it has about the underlying system [8]. In standard QEC schemes such as surface code quantum computing [9, 4], successive rounds of stabiliser measurements are performed to track the state of a logical qubit against potential errors. This measurement data is typically converted from an analogue to a digital representation and classified to a binary measurement outcome before being passed to a classical decoder [10]. However, the binary classification process is sub-optimal as it throws away valuable information about the measurement process that can be used to improve decoding accuracy [11, 12]. By using a probability estimator that captures the likelihood of a given measurement signal originating from state |$\ket{0}$ or |$\ket{1}$, denoted as soft information, the decoder can more accurately track measurement errors over the course of a QEC protocol, enhancing the code threshold and making the decoder more resilient to classification errors.
Soft information has been shown to provide lower logical error rates [8, 13, 14, 15], higher code thresholds [12] and stronger error suppression than conventional hard decoding [16]. However, the impact of soft information on the sub-threshold scaling and physical qubit footprint of error-corrected systems has been under-explored. In this work, we investigate this. We find that decoding using soft information, which we call soft decoding, gives consistently higher rates of error suppression than hard decoding, with the difference between the two growing as a function of the amount of measurement classification errors in the system.
To evaluate the impact of soft decoding on the scaling properties of error-corrected platforms, we simulate two leading qubit candidates for large-scale fault-tolerant quantum computing—superconducting and neutral-atom qubits. While superconducting qubits have enjoyed significant successes towards scalable QEC, neutral atoms have emerged as a competitive alternative thanks to their long coherence times, high-fidelity scalable gates, long-range connectivity and simultaneous manipulation of multi-qubit arrays [17, 2, 18]. On both platforms, we evaluate the rate of error suppression, represented by the $\Lambda$-factor [19], that can be achieved with soft information decoding and compare it to traditional hard decoding. To quantify error suppression, we simulate quantum memories encoded in surface codes and syndrome extraction circuits in the recently proposed bivariate bicycle (BB) codes [20]. We use a soft-information-aware variant of the local clustering decoder (LCD) [21] to decode the surface code simulations, and a belief propagation (BP) based decoder [22, 23] augmented with soft information to decode the BB codes.
@@
We see that in simulated quantum memory experiments, soft decoders provide consistently higher $\Lambda$-factors than hard decoders for both superconducting and neutral atom platforms. In noise regimes where measurement classification errors are unlikely compared to data errors, soft decoding offers next to no advantage, but in more realistic settings where measurement classification is a lower-fidelity operation than gates, soft decoding offers between $10\%$- and $20\%$-improvements in $\Lambda$. Our modelling shows that the scaling advantages of soft decoding are maintained even when measurement times are shorter than ideal—offering faster QEC cycles without a penalty in the error suppression performance compared to hard decoders. For simulated bivariate bicycle codes, we find soft information to reduce logical error probabilities by an order of magnitude, showcasing that the benefits of the technique are applicable across a wide range of QEC codes.
@@
The paper is structured as follows. A summary of the soft measurement model for superconducting and neutral atom platforms is described in section II.1, and our chosen figures of merit are described in section II.2. In section II.3, we study the impact of soft decoding for surface code experiments on both superconducting and neutral atom platforms, leveraging soft and hard variants of LCD as our decoder. In section II.3.2, we simulate error suppression against the readout duration of superconducting and neutral atom platforms, likewise decoding the results with soft and hard LCD. In section II.4 we show how soft-BP decoders benefit from soft measurement information, benchmarking the decoders on BB codes on simulated neutral atom and superconducting platforms. Section III concludes the work and provides pointers on effectively incorporating soft information into the decoding workflows of future QEC experiments.
II Results
@@ II.1 Soft measurement model
When measuring the state of a qubit, we treat the resulting signal as a random variable $\mu$ which follows some probability density function $f^{(\bar{\mu}=0)}(\mu)$ for a qubit projected into the |$\ket{0}$-state, and a density function $f^{(\bar{\mu}=1)}(\mu)$ for a qubit projected into the |$\ket{1}$-state. The type of the measurement distribution depends on the physical mechanism of the measurement process, determined by the qubit type [24, 25]. We label the state to which the qubit was projected by the variable $\bar{\mu}$, here referred as the ideal outcome, where $\bar{\mu}=0$ and $\bar{\mu}=1$ correspond to states |$\ket{0}$ and |$\ket{1}$ respectively. To classify a soft measurement response $\mu$ to a binary outcome $\hat{\mu}\in\{0,1\}$, we calculate from each measurement the posterior probability
@@
Figure 1: Readout model for superconducting and neutral atom qubits. I. Superconducting readout, a) showing example IQ voltages from a repeated state-preparation and measurement routine for states |$\ket{0}$ (in purple) and |$\ket{1}$ (in orange), b) the resulting probability density functions when projected along the axis connecting the 0- and 1-centroids, and c) the posterior probability of a measurement outcome $\mu$ given the initial state preparation |$\ket{0}$ or |$\ket{1}$. II. Neutral atom readout, a) showing photon counts $\mu$ for a bright state |$\ket{1}$ (in orange) and a dark state |$\ket{0}$ (in purple), and b) the posterior probability of a measurement outcome $\mu$ given either a 0- or 1-state preparation.
The classification outcome for a soft measurement $\mu$, also denoted the hardened measurement outcome, is labelled by $\hat{\mu}$. Given a posterior probability $P(1\mid\mu)<1/2$, the classification outcome is $\hat{\mu}=0$, while $P(1\mid\mu)\geq 1/2$ results in a classification outcome $\hat{\mu}=1$. To simplify the expression Eq. 1, we assume for our simulations that $P(0)=P(1)=1/2$, which is the case for long QEC experiments where the outcomes of ancilla qubit measurements are unbiased [16]. While the physical origin of the measurement probability distribution differs based on the hardware type and can vary qubit-to-qubit, the posterior probability $P(1\mid\mu)$ can be easily mapped to a binary measurement outcome $\hat{\mu}\in\{0,1\}$ and processed without prior knowledge of the measurement model, making it a broadly applicable input to different decoder types, such as matching-based [26] or neural-network decoders [8].
In the case of superconducting qubits, shown in Fig. 1 I) the measurement signal comes as a dispersive shift of a resonator that is coupled to the qubit, resulting in a unique Gaussian distribution depending on the qubit state [25]. For neutral atoms, shown in Fig. 1 II) we model a non-destructive fluorescence-based state-readout protocol that is compatible with mid-circuit measurements. Measurement is performed using a single photon counting module (SPCM), with the |$\ket{0}$- and |$\ket{1}$-states exhibiting different amounts of fluorescence [27, 28]. The resulting measurement signal is a unique Poisson distribution of the photon count $\mu$ for the two states [29]. The probability density functions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$ used for our simulations of superconducting and neutral atom qubits are given in section IV.1.1 and section IV.1.2 respectively. In both platforms, there exists a classification boundary at a fixed value of $\mu$ where the measurement outcome is maximally uncertain and classification errors are most likely. On either side of the boundary, measurement outcomes are more likely to be correctly assigned.
For both platforms, the choice of measurement time is a key parameter affecting the measurement fidelity—we demonstrate its effect in detail in section II.3.2. In order to mitigate the random fluctuations inherent in the measurement process, measurement signal is accumulated over a fixed time interval, with long measurement durations resulting in a cleaner signal. However, even with a good choice of measurement time, it is not possible to always distinguish between the two possible states, and cases where an incorrect state is assigned are known as soft measurement errors or soft flips [12]. Given a soft measurement value $\mu$, the probability of a soft flip is given by
@@
When binary measurements are used, the duration of the measurement operation is typically optimised for maximum measurement fidelity $1-p_{\textrm{S}}$ [30, 31], but simulations show that a shorter readout duration can result in a higher logical fidelity when QEC is applied thanks to a reduction in idling errors during measurement [12]. In our model, used in section II.3 and section II.4, we use measurement times $\tau_{\textrm{M}}$ and error probabilities according to current state-of-the art devices from Google Quantum AI [1] for superconducting devices, and a $Z$-biased noise model along with an appropriate measurement model for neutral atom based devices [32, 24].
@@ II.2 Figures of merit
Our metric of interest is the error suppression rate $\Lambda$, which quantifies how much the logical error rate of a given QEC code decreases when the code distance is increased [19]. The distance, denoted $d$, defines how many errors are required for the code to experience an undetectable logical failure, with bigger distances making the code more resilient to errors. The logical error rate $\epsilon_{\textrm{L}}$ of a QEC code with distance $d$ is suppressed according to
$\epsilon_{\textrm{L}}\propto\Lambda^{-(d+1)/2}\quad.$ (4)
@@
To evaluate $\Lambda$, we generate instances of rotated planar codes with varying code distances $d=\{5,7,9\}$ and simulate the logical fidelity of a quantum memory experiment taking $T$ syndrome extraction rounds. By decoding the resulting syndrome measurement data, we compare the state of the logical qubit at the end of the experiment with decoder corrections applied against the known state of the logical qubit. By repeating the experiment over multiple shots, we obtain the logical error probability $P_{\textrm{L}}(d,T)$ for each code distance $d$ and number of rounds $T$. To extract $\Lambda$ from this data, we first use the relation
$P_{\textrm{L}}(d,T)=\frac{1}{2}\left[1-(1-2\epsilon_{\textrm{L}}(d))^{T}\right]$ (5)
to compute the logical error rate per round $\epsilon_{\textrm{L}}(d)$, and then use a linear fit of $\log(\epsilon_{\textrm{L}})$ to compute $\Lambda$ the details of this process are given in section IV.2.
@@
NA $p$ $0.5\%$ $(6\pm 1)$ $(7\pm 2)$ $(16\pm 5)\%$
NA $p$ $1\%$ $(2.9\pm 0.2)$ $(3.0\pm 0.2)$ $(3.0\pm 0.2)\%$
NA $5p$ $0.5\%$ $(2.2\pm 0.1)$ $(2.7\pm 0.2)$ $(20\pm 2)\%$
NA $5p$ $1\%$ $(1.26\pm 0.02)$ $(1.44\pm 0.03)$ $(14.6\pm 0.3)\%$
SC $p$ $0.3\%$ $(1.86\pm 0.03)$ $(1.90\pm 0.03)$ $(2.3\pm 0.1)\%$
SC $p$ $0.5\%$ $(1.18\pm 0.01)$ $(1.21\pm 0.01)$ $(2.0\pm 0.2)\%$
SC $5p$ $0.3\%$ $(1.61\pm 0.02)$ $(1.78\pm 0.02)$ $(10.4\pm 0.2)\%$
SC $5p$ $0.5\%$ $(1.07\pm 0.01)$ $(1.13\pm 0.01)$ $(6.4\pm 0.1)\%$
Table 1: Performance of soft decoders on simulated quantum memory experiments on superconducting (SC) and neutral atom (NA) platforms. For neutral atom platforms, we use a $Z$-biased noise model with bias factor $100$. Both the neutral atom and the superconducting simulations use circuit-level noise with error rate $p$, as described in section IV.1.2 and section IV.1.1 respectively. We vary the amount of classification error $p_{\textrm{S}}$ in the system, and show the average error suppression factor $\Lambda$ for soft and hard variants of the local clustering decoder. Each experiment uses $T=10$ syndrome extraction rounds, and is repeated for $10^{6}$ shots.
II.3 Soft decoding with LCD
II.3.1 Quantum memory performance
To evaluate decoding accuracy, we simulate quantum memory experiments for superconducting and neutral atom qubits under a circuit-level noise model using the fast Clifford simulator Stim [33]. Our noise models have two parameters—the two-qubit gate depolarising error probability $p$ and the average soft measurement error probability $p_{\textrm{S}}$. In the case of superconducting qubits, we use a variant of the SI1000 model [34] to parametrise the noise channels in our simulations, described in detail in section IV.1.1. For our neutral atom simulations we use a $Z$-biased error model with a bias factor of $100$, leading to $Z$-errors with probability $p/3$ and $X$- and $Y$-errors with probability $p/300$. The full noise model for neutral atoms is given in section IV.1.2. For both platforms, we introduce soft information to the simulation by evaluating the posterior probabilities $P(1\mid\mu)$ as defined in Eq. 1. The ideal outcome $j$ used to draw a sample from the platform-specific measurement distributions $f^{(\bar{\mu}=j)}(\mu)$ is determined by the Stim measurement sample, and the posterior probability is evaluated accordingly.

### v3

@@ Abstract
Imperfect measurements are a prevalent source of error across quantum computing platforms, significantly degrading the logical error rates achievable on current hardware. To mitigate this issue, rich measurement data referred to as soft information has been proposed to efficiently identify and correct measurement errors as they occur. In this work, we model soft information decoding across a variety of physical qubit platforms and decoders and showcase how soft information can make error correction viable at lower code distances and higher physical error rates than is otherwise possible. We simulate the effects of soft information decoding on quantum memories for surface codes and bivariate bicycle codes, and evaluate the error suppression performance of soft decoders against traditional decoders. Our simulations show that soft information decoding on near-term devices can provide up to $11\%$ higher error suppression on superconducting qubits and up to $20\%$ stronger error suppression on neutral atom qubits. These accuracy gains correspond to $13\%$ and $33\%$ reductions in the physical qubit footprint of superconducting and neutral atom devices respectively when operating at a logical error rate of $10^{-6}$, showcasing that soft information is a powerful tool for reducing the cost and complexity of large-scale fault-tolerant quantum computers.
@@ I Introduction
[…]
Prior simulations have shown that soft information can achieve a higher error-correction threshold on the surface code than traditional protocols [12], and a number of experimental works have confirmed that the technique yields lower logical error rates in real-world systems [13, 14, 15, 16]. The primary question that had been left to answer is whether soft information remains advantageous when moving towards large-scale future systems, and whether the previously shown improvements translate to other qubit platforms and error-correcting codes. It also remained an open question whether soft decoding is compatible with the high-throughput, low-latency decoding workflows needed for future large-scale error-corrected systems [1]. Previous soft decoder demonstrations required extensive computation during runtime [12, 16], which are potentially incompatible with real-time implementations.
To fill this gap, we incorporate soft decoding capabilities into two promising real-time decoder architectures — the local clustering decoder (LCD) [17] and belief propagation (BP) [18, 19, 20], and benchmark their accuracy on extensive simulations of two leading qubit platforms: superconducting transmon qubits and neutral atom qubits. Although superconducting qubits have enjoyed significant successes towards scalable QEC, neutral atoms have emerged as a competitive alternative thanks to their high fidelity, extensive connectivity and the ability to simultaneously manipulate multi-qubit arrays [21, 2, 22]. On these platforms, we simulate logical performance of our decoders on the surface code as well as the newly developed bivariate bicycle (BB) codes [23], and compare the resource overheads needed to reach target error rates with soft and hard decoders. Our findings show that soft information yields consistently higher rates of error suppression than hard decoding across different error-correcting codes and qubit platforms, unlocking low logical error rates at substantially smaller qubit footprints that would otherwise be needed. Our results are of particular relevance to experimentalists, who we hope will leverage soft information to design faster and more hardware-efficient error-corrected quantum devices in the future.
For our soft-information-augmented decoders, we use LCD [17] to decode the surface code simulations, and the BP-based decoder [18, 19] to decode the BB codes. Our choice of decoders is motivated by their prospects in a real-time experiment — LCD is optimised for high-throughput implementation, and BP variants such as relay [20] have been shown to be compatible with the tight time constraints of real-time decoding. While most small-scale QEC experiments have been conducted with offline decoders where the syndrome information is decoded after the fact, utility-scale quantum computers will require data processing and decoding to happen at the timescale of syndrome extraction — requiring scalable and responsive decoders. We note that other decoders can be made compatible with soft information, such as minimum-weight perfect matching [24] which has been shown to deliver improved error rates in real-world experiments [13, 14, 16], as well as network decoders which can be trained with soft information, delivering better error rates than their binary counterparts [8, 25]. More general solvers such as Tesseract [26] may also be used with soft information by modifying the decoder prior, but these lack the scalability needed for a real-time setting.
@@
On both platforms, we evaluate the rate of error suppression, represented by the $\Lambda$-factor [27], that can be achieved with soft information decoding and compare it to traditional hard decoding. We then relate improvements in $\Lambda$ to physical qubit footprint estimates given target logical error rates. We see that in simulated quantum memory experiments, soft decoders provide consistently higher $\Lambda$-factors than hard decoders for both superconducting and neutral atom platforms. In noise regimes where measurement classification errors are unlikely compared to data errors, soft decoding offers next to no advantage, but in more realistic settings where measurement classification is a lower-fidelity operation than gates, soft decoding offers between $11\%$- and $20\%$-improvements in $\Lambda$. This corresponds to $13\%$ and $33\%$ reductions in physical qubits needed to reach a logical error rate of $10^{-6}$ for superconducting and neutral atom platforms respectively. Our modelling shows that the scaling advantages of soft decoding are maintained even when measurement times are shorter than ideal—offering faster QEC cycles without a penalty in the error suppression performance compared to hard decoders. For simulated bivariate bicycle codes, we find soft information to reduce logical error probabilities by an order of magnitude, showcasing that the benefits of the technique are applicable across a wide range of QEC codes.
@@
The paper is structured as follows. A summary of the soft measurement model for superconducting and neutral atom platforms is described in section II.1, and our chosen figures of merit are described in section II.2. In section III.1, we study the impact of soft decoding for surface code experiments on both superconducting and neutral atom platforms, leveraging soft and hard variants of LCD as our decoder. In section III.1.2, we simulate error suppression against the readout duration of superconducting and neutral atom platforms, likewise decoding the results with soft and hard LCD. In section III.2 we show how soft-BP decoders benefit from soft measurement information, benchmarking the decoders on BB codes on simulated neutral atom and superconducting platforms. Section IV concludes the work and provides pointers on effectively incorporating soft information into the decoding workflows of future QEC experiments.
II Method
@@ II.1 Soft measurement model
[…]
@@
[…]
For both platforms, the choice of measurement time is a key parameter affecting the measurement fidelity—we demonstrate its effect in detail in section III.1.2. In order to mitigate the random fluctuations inherent in the measurement process, measurement signal is accumulated over a fixed time interval, with long measurement durations resulting in a cleaner signal. However, even with a good choice of measurement time, it is not possible to always distinguish between the two possible states, and cases where an incorrect state is assigned are known as soft measurement errors or soft flips [12]. Given a soft measurement value $\mu$, the probability of a soft flip is given by
@@
When binary measurements are used, the duration of the measurement operation is typically optimised for maximum measurement fidelity $1-p_{\textrm{S}}$ [33, 34], but simulations show that a shorter readout duration can result in a higher logical fidelity when QEC is applied thanks to a reduction in idling errors during measurement [12]. In our model, used in section III.1 and section III.2, we use measurement times $\tau_{\textrm{M}}$ and error probabilities according to current state-of-the art devices from Google Quantum AI [1] for superconducting devices, and a $Z$-biased noise model along with an appropriate measurement model for neutral atom based devices [35, 28].
It is crucial to note that soft measurement flips are a distinct error mechanism compared to bit-flip errors that occur on ancilla qubits prior to measurement. If ancilla qubits are not reset between syndrome extraction rounds, the error $e_{\textrm{S}}$ from a soft measurement flip shows up in two fault locations separated by two time steps [13, 36], whereas a bit-flip error $e_{\textrm{B}}$ triggers a pair of fault locations separated by a single time step. This difference stems from the fact that a soft measurement error does not affect the true state of the ancilla qubit, while a bit-flip error does. Given a soft measurement flip channel with probability $p_{\textrm{S}}$ and a measurement bit-flip channel with probability $p_{\textrm{B}}$, a QEC circuit without resets will result in a decoding graph with two distinct edges $e_{\textrm{S}}\neq e_{\textrm{B}}$. In the presence of resets, the two edges have the same endpoints and the measurement error channel can be collapsed into a single probability $p_{\textrm{M}}=p_{\textrm{S}}(1-p_{\textrm{B}})+p_{\textrm{B}}(1-p_{\textrm{S}})$. The distinct nature of soft measurement errors compared to other measurement error mechanisms plays an important role in designing optimal syndrome extraction circuits  [36].
@@ II.2 Figures of merit
Our metric of interest is the error suppression rate $\Lambda$, which quantifies how much the logical error rate of a given QEC code decreases when the code distance is increased [27]. The distance, denoted $d$, defines how many errors are required for the code to experience an undetectable logical failure, with bigger distances making the code more resilient to errors. The logical error rate $\epsilon_{d}$ per QEC round of a code with distance $d$ is suppressed according to
$\epsilon_{d}\propto\Lambda^{-\frac{d+1}{2}}\quad.$ (4)
@@
To evaluate $\Lambda$, we generate instances of rotated planar codes with varying code distances $d=\{5,7,9\}$ and simulate the logical fidelity of a quantum memory experiment taking a fixed number $T=10$ syndrome extraction rounds. By decoding the resulting syndrome measurement data, we compare the state of the logical qubit at the end of the experiment with decoder corrections applied against the known state of the logical qubit. By repeating the experiment over multiple shots, we obtain the logical error probability $P_{\textrm{L}}(T)$ for each code distance $d$ and number of rounds $T$. To extract $\Lambda$ from this data, we first use the relation
$P_{\textrm{L}}(T)=\frac{1}{2}\left[1-(1-2\epsilon_{d})^{T}\right]$ (5)
to compute the logical error rate per round $\epsilon_{d}$ [1], and then use a linear fit of $\log(\epsilon_{d})$ to compute $\Lambda$ — the details of this process are given in section VI.2. For consistency, we always use a fitted $\Lambda$ as opposed to a distance-specific value and omit the $d=3$ data point due to finite-size effects [37].
Given some $p_{0}$ and $\Lambda$ as obtained from a curve fit, we evaluate physical qubit footprints required for a given target error rate by solving Eq. 5 for $d$ given a fixed $P_{\textrm{L}}$, assuming $T=d$ syndrome extraction rounds  [^1] This assumes lattice surgery as the means of logical computation. For qubit platforms such as neutral atoms that are compatible with transversal logic, only $\mathcal{O}(1)$ rounds are needed for logical operations and one can achieve the same logical error rate with a lower-distance code. Using the definition of a quantum operation (Quop) [39], we pick two key logical error rate regimes as benchmarks for device performance:
  • KiloQuop: $P_{L}=10^{-3}$, corresponding to $\sim 10^{3}$ error-free operations. This is an error regime that near-future machines are likely to achieve.
  • MegaQuop: $P_{L}=10^{-6}$, corresponding to $\sim 10^{6}$ error-free operations. This error regime is much harder to attain, and sets the benchmark for future devices.
Given a target regime from the above, we estimate the physical qubit footprint $N$ of a rotated planar surface code for each decoder according to $N=2d_{\textrm{min}}^{2}-1$ where $d_{\textrm{min}}$ is the smallest odd distance that satisfies a logical error rate per $d$ rounds of QEC lower than $10^{-3}$ or $10^{-6}$ for the KiloQuop and MegaQuop regimes, respectively. Our estimates are based on the fitted $\Lambda$, meaning that footprint values that require $d>9$ are extrapolations from the fit.
@@
III Results
III.1 Soft decoding with LCD
III.1.1 Quantum memory performance
To evaluate decoding accuracy, we simulate quantum memory experiments for superconducting and neutral atom qubits under a circuit-level noise model using the fast Clifford simulator Stim [40]. Our noise models have two parameters—the two-qubit gate depolarising error probability $p$ and the average soft measurement error probability $p_{\textrm{S}}$. In the case of superconducting qubits, we use a variant of the SI1000 model [41] to parametrise the noise channels in our simulations, described in detail in section VI.1.1. We evaluate logical performance under physical error rate regimes between $p=0.2\%$ and $p=0.7\%$, reflecting recent real-world benchmarks of transmon qubits operating near the error-correction threshold [1, 42]. The threshold (estimated at $p=0.6\%$ in Ref. [12]) provides an upper bound on the effectiveness of soft decoding for the surface code, so we focus our analysis on the sub-threshold scaling behaviour. For our neutral atom simulations we use a $Z$-biased error model with a bias factor of $100$, leading to $Z$-errors with probability $p/3$ and $X$- and $Y$-errors with probability $p/300$. The full noise model for neutral atoms is given in section VI.1.2. Here, we sweep physical error rates from $p=0.5\%$ to $p=1\%$, reflecting the two-qubit gate error rates of recent experiments on neutral atom qubits [43, 21].
We leverage Stim to generate ideal measurement outcomes $j$ that include all noise channels except for measurement classification errors and pass these to the platform-specific measurement distributions $f^{(\bar{\mu}=j)}(\mu)$ to sample soft measurement values $\mu$. Given the soft measurement samples we evaluate the posterior probability $P(1\mid\mu)$ via Eq. 1, which is fed to our decoder along with the syndrome extracted from the hardened measurement outcomes $\hat{\mu}$.